# online_appendix

### v1

@@ 3 What explains the increase in the number of low-value trade transactions?
The fast-fashion business model requires stores to replenish frequently, a procedure that echoes the just-in-time system, introduced first by Toyota in the automotive sector and later implemented in other industries (Womack et al.,, 1990). Since all products exported by leading Spanish retailers go through a distribution center in Spain (Ghemawat and Nueno,, 2006; Rodríguez,, 2024), we expect the fast fashion strategy to generate a large number of low-value transactions between the distribution center and stores abroad. [^12] We define as leading Spanish retailers all the brands owned by the top 2 Spanish fashion groups. We also anticipate the frequency of transactions to be higher in destinations geographically far from Spain. Because products take longer to reach their destination, firms have to respond very quickly to replenishment demands from stores to ensure that preferences do not change before goods arrive. This will demand the use of air transport to serve the destinations that cannot be reach by road in few days from Spain.
@@
Using the procedure described in de Lucio et al., (2018), we identified the export transactions of the leading fashion retailers in Spain in our database. Panel A of Figure 3 shows the evolution of the number of low-value export transactions of the leading fashion retailers and all firms between 1997 and 2023. From 1997 to 2021, the increase in the number of low-value export transactions for all firms runs parallel to the increase in the number of low-value export transactions from the leading fashion retailers. This is because exports from the leading fashion retailers represented most of the volume of low-value export transactions. For example, in 2021, the leading fashion retailers represented 63% of the total number of low-value export transactions. Panel B of Figure 3 shows that the top 10 destinations of the low-value export transactions of the leading fashion retailers in 2021 were countries that were not geographically close to Spain. [^13] The top 10 destinations accounted for 55% of the total number of low-value export transactions. In fact, countries that are closer to Spain, such as France, Germany, Italy, or Portugal, and which were some of the most important export destinations by value of the leading fashion retailers in 2021, were not included in the top 10 ranking. Furthermore, except for destinations that could be reached by road in a few days from Spain (Russian Federation, Turkey, and the UK), most of the products were transported by air. These findings are in line with the fast-fashion strategy described above. In summary, from 1997 to 2021, the leading fashion retailers were responsible for the increase in the total number of low-value export transactions. The frequency of low-value transactions, their geographical distribution, and the means used to transport the products are consistent with the fast-fashion strategy followed by these firms.
@@
The number of low-value export transactions of the leading Spanish retailers decreases from 2021 onward. A factor contributing to this decline was the termination or reduction of operations in Russia after the invasion of Ukraine in February 2022 (de Lucio et al.,, 2024). [^14] One million export transactions to Russia from the leading fashion retailers were lost after the invasion of Ukraine. However, the low-value export transactions of the leading fashion retailers continued in decrease in 2023, suggesting that other reasons were also explaining this decline. Furthermore, the total number of low-value export transactions decreases less than the number of low-value export transactions of the leading fashion retailers. This indicates that firms not related to the fashion industry began to account for a sizable share of low-value export transactions. To explain this fact, we need to introduce the second explanation for the rise in the number of low-value transactions. We do it analyzing, first, the evolution of the number of low-value import transactions.
@@
B. Top 10 destinations for low-value export transactions from leading fashion retailers, 2021
@@ C. Number and share of low-value transactions in Spain’s imports from China, 1997-2023
Our data set does not provide information to determine whether a transaction is the result of an online or an offline purchase. Having this limitation in mind, we examine whether other facts are consistent with the direct-to-consumer explanation for the rise in the number of low-value import transactions from China from 2020 onward. Our data set shows that in 2023, only one firm, which we cannot identify by name, concentrated almost half of low-value import transactions originating in China. This firm did not receive imports from China in a specific province in Spain, but they were scattered throughout all Spanish provinces. Furthermore, this firm recorded 300,000 different import transactions from China in each monthly customs declaration. This very large number of declared transactions is consistent with the fact that online platforms, or the logistic partners of online platforms, can use a simplified version of the customs declaration to record transactions that have a value equal to or below the 150 euros de minimis threshold (the H7 declaration). In this simplified declaration, the operator should record every low-value import shipment destined to each customer. It is very unlikely that a firm would fraction imports for its own use into so many low-value transactions to avoid the payment of tariffs, because such a large number of transactions would raise the suspicion of tax authorities. These facts (the geographical distribution and the large number of transactions) are consistent with the argument that products were sent directly from Chinese producers to Spanish customers and that the importer was a the operator that facilitated customs clearance and shipments.
@@
The top importer by number of low-value transactions from China in 2023 was the same as in 2022. However, before 2022, the main importers were a large retailer and two fashion firms that we were able to identify by name. [^16] For confidentiality issues, we cannot provide the name of these firms. We also cannot offer the references that back our assertions because they would enable the identification of the firms. These references can be requested from the authors. These firms received all low-value imports from China in the province where their headquarters and main warehouse were located. We do not have information to determine whether the increase in the number of low-value import transactions from China of these firms was explained by the expansion of their e-commerce sales or the use of a just-in-time strategy. However, there are two facts that are consistent with the argument that these purchases were likely to be linked to these firms’ e-commerce activities.
@@
First, in the case of the large retailer, the number of low-value import transactions from China was almost negligible in 2019. However, it became the importer with the largest number of low-value transactions from China in 2020. That year, due to Covid, non-essential stores remained closed for some weeks. Furthermore, even when the stores opened after the strictest confinement measures were lifted, the shopping experience became less pleasant for the customers due to health-related restrictions. In parallel, the retailer reported a very large increase in e-commerce sales between 2019 and 2020. These facts suggest that it was the increase in purchases on this firm’s online marketplace which contributed to the increase in the number of low-value import transactions from China. Second, as argued above, the fact that the large retailer and the two fashion firms reported very large numbers of low-value import transactions in their customs declarations is consistent with the documentation of e-commerce-based transactions.
@@
We find that the four firms that explain the rise in the number of low-value import transactions from China are also responsible for one third of the total increase in the number of low-value import transactions between 1997 and 2023. The remaining 100,000 firms that performed low-value import transactions each contributed a very tiny fraction to the remaining two-thirds of the increase. [^17] The exceptions are two firms that represented 4% and 2% of the increase that we could not identify. Both firms specialize in low-value imports from Germany.
@@
At this point, we address the increase in the number of low-value export transactions that occurred between 2021 and 2023. The firm that contributes the most to the growth of low-value transactions in this period (16% of the total) is a retailer that we were able to identify by name. The increase in the number of low-value export transactions coincides with the opening of online marketplaces in three European countries by this firm since 2021. These marketplaces allowed Spanish firms that supplied the retailer in Spain to sell their products on other European markets. Consistent with this narrative, we observe a large increase in the number of low-value exports by the retailer from 2021 onward, particularly in those European countries in which the marketplace was launched. The emergence of this retailer as a major player in the volume of export transactions also explains why there are differences in the ranking of destinations for low-value export transactions between 2021 and 2023. In 2021, low-value export transaction destinations were still dominated by the destinations of leading fashion retailers (panel B of Figure 3). However, by 2023, export destinations were also affected by the new retailer’s destinations, which were concentrated in European countries geographically close to Spain.
@@
Taking into account the caveats of our data set, which cannot discriminate between online and offline transactions, these results suggest that online marketplaces could also play a relevant role in explaining the rise in the number of low-value transactions in exports. They also underscore how decisions taken by few firms can have very large effects on the number of low-value trade transactions.
@@ 4 Conclusion
This paper has documented a novel feature of international trade: the rise in the number of low-value transactions. Using the universe of Spanish international trade transactions, we have shown that the share of low-value transactions in the total number of transactions between 1997 and 2023 increased from 9% to 61% in exports and from 14% to 54% in imports.
We have offered two explanations for the increase in the number of low-value trade transactions. First, the rise in the number of low-value export transactions, at least until 2021, is explained by transactions by fashion retailers. This high volume of low-value transactions is consistent with a fast-fashion strategy, which uses a frequent stocking of stores to adjust for changes in consumer preferences and creates a sense of scarcity among their customers. Second, the increase in the number of low-value transactions is consistent with the emergence of online retail platforms that allow consumers to have easy access to a wide range of low-value items manufactured abroad.
@@
Figure A.1: Robustness analyses on the share of low-value transactions in total number of trade transactions
@@ A. Alternative low-value transaction thresholds
Note: Real de minimis corresponds to the 150 euro threshold, established in 2009, adjusted by the Spanish consumer price index (Spanish Statistical Institute, available at https://www.ine.es). US de minimis is the de minimis 800 USD value established by the United States in 2016. Benchmark corresponds to the share of low-value transactions reported in Panel A of Figure 1 in the main text.
@@ B. Top 20 countries by share in the total number of high-value transactions
Figure A.4: Ranking of number of low-value transactions by destinations and origins in 2021

### v2

@@ 3 What explains the increase in the number of low-value trade transactions?
The fast-fashion business model requires stores to replenish frequently, a procedure that echoes the just-in-time system, introduced first by Toyota in the automotive sector and later implemented in other industries (Womack et al.,, 1990). Since all products exported by leading Spanish retailers go through a distribution center in Spain (Ghemawat and Nueno,, 2006; Rodríguez,, 2024), we expect the fast fashion strategy to generate a large number of low-value export transactions between the distribution center and stores abroad. [^16] We define as leading Spanish retailers all the brands owned by the top 2 Spanish fashion groups. We also anticipate the frequency of transactions to be higher in destinations geographically far from Spain. Because products take longer to reach their destination, firms have to respond very quickly to replenishment demands from stores to ensure that preferences do not change before goods arrive. This will require the use of air transport to serve destinations that cannot be reached by road in a few days from Spain.
@@
Using the procedure described in de Lucio et al., (2018), we identified the export transactions of the leading fashion retailers in Spain in our database. Panel A of Figure 3 shows the evolution of the number of low-value export transactions of the leading fashion retailers and all firms between 1997 and 2023. From 1997 to 2021, the increase in the number of low-value export transactions for all firms runs parallel to the increase in the number of low-value export transactions from the leading fashion retailers. This is because exports from the leading fashion retailers represented most of the volume of low-value export transactions: 63% in 2021. Panel B of Figure 3 shows that the top 10 destinations of the low-value export transactions of the leading fashion retailers in 2021 were countries that were not geographically close to Spain. [^17] The top 10 destinations accounted for 55% of the total number of low-value export transactions. In fact, countries that are closer to Spain, such as France, Germany, Italy, or Portugal, and which were some of the most important export destinations by value of the leading fashion retailers in 2021, were not included in the top 10 ranking. Furthermore, except for destinations that could be reached by road in a few days from Spain (Russian Federation, Turkey, and the UK), most of the products were transported by air. These findings are in line with the fast-fashion strategy described above. In summary, from 1997 to 2021, the leading fashion retailers were responsible for the increase in the total number of low-value export transactions. The frequency of low-value transactions, their geographical distribution, and the means used to transport the products are consistent with the fast-fashion strategy followed by these firms.
@@
The number of low-value export transactions of the leading Spanish retailers decreases from 2021 onward. A factor contributing to this decline was the termination or reduction of operations in Russia after the invasion of Ukraine in February 2022 (de Lucio et al.,, 2024). [^18] One million export transactions to Russia from the leading fashion retailers were lost after the invasion of Ukraine. The total number of low-value export transactions decreases less than the number of low-value export transactions of the leading fashion retailers. This indicates that firms not related to the fashion industry began to account for a larger share of low-value export transactions. To explain this fact, we need to introduce the second explanation for the increase in the number of low-value transactions. We do it analyzing, first, the evolution of the number of low-value import transactions.
@@
B. Top 10 destinations by volume for low-value export transactions from leading fashion retailers, 2021
@@ C. Number and share of low-value transactions in Spain’s imports from China, 1997-2023
Our data set does not provide information to determine whether a transaction is the result of an online or an offline purchase. Having this limitation in mind, we examine whether other facts are consistent with the direct-to-consumer explanation for the rise in the number of low-value import transactions from China from 2020 onward. In 2023, a firm concentrated almost half of low-value import transactions originating in China. This firm did not receive imports from China in a specific province in Spain, but they were scattered throughout all Spanish provinces. Furthermore, this firm recorded 300,000 different import transactions from China in each monthly customs declaration. This very large number of declared transactions is consistent with the fact that online platforms, or their logistic partners, can fulfill the customs declaration on behalf of their final customers, when the transaction is equal or below the EU de minimis threshold. It is very unlikely that a firm would fraction imports for its own use into so many low-value transactions to avoid the payment of tariffs, because such a large number of transactions would raise the suspicion of tax authorities. Therefore, the geographical distribution and the large number of transactions are consistent with the argument that products were sent directly from Chinese producers to Spanish customers and that the importer was an operator that facilitated customs clearance and shipments.
@@
The top importer by number of low-value transactions from China in 2023 was the same as in 2022. However, before 2022, the main importers were a large retailer and two fashion firms. They firms received all low-value imports from China in the province where their headquarters and main warehouse were located. We do not have information to determine whether the increase in the number of low-value import transactions from China of these firms was explained by the expansion of their e-commerce sales or the use of a just-in-time strategy. However, there are two facts that are consistent with the argument that these purchases were likely to be linked to these firms’ e-commerce activities.
@@
First, in the case of the large retailer, the number of low-value import transactions from China was almost negligible in 2019. However, it became the importer with the largest number of low-value transactions from China in 2020. That year, due to Covid, non-essential stores remained closed for some weeks. Furthermore, even when the stores opened after the strictest confinement measures were lifted, the shopping experience became less pleasant for the customers due to health-related restrictions. In parallel, the retailer reported a very large increase in e-commerce sales between 2019 and 2020. These facts suggest that it was the increase in purchases on this firm’s online marketplace that contributed to the increase in the number of low-value import transactions from China. Second, as argued above, the fact that the large retailer and the two fashion firms reported very large numbers of low-value import transactions in their customs declarations is consistent with the documentation of e-commerce-based transactions.
@@
We find that the four firms that explain the rise in the number of low-value import transactions from China are also responsible for one third of the total increase in the number of low-value import transactions between 1997 and 2023. The remaining 100,000 firms that performed low-value import transactions each contributed a very tiny fraction to the remaining two-thirds of the increase. [^20] The exceptions are two firms that represented 4% and 2% of the increase that we could not identify by name. Both firms specialize in low-value imports from Germany.
@@
At this point, we address the increase in the number of low-value export transactions that occurred between 2021 and 2023. The firm that contributes the most to the growth of low-value transactions in this period (16% of the total) is a retailer. The increase in the number of low-value export transactions coincides with the opening of online marketplaces in three European countries by this firm since 2021. These marketplaces allowed Spanish firms that supplied the retailer in Spain to sell their products on other European markets. Consistent with this narrative, we observe a large increase in the number of low-value export transactions by the retailer from 2021 onward, particularly in those European countries in which a marketplace was launched. The emergence of this retailer as a major player in the volume of export transactions also explains why there are differences in the ranking of destinations for low-value export transactions between 2021 and 2023. In 2021, low-value export transaction destinations were still dominated by the destinations of leading fashion retailers (panel B of Figure 3). However, by 2023, export destinations were also affected by the new retailer’s destinations, which were concentrated in European countries geographically close to Spain.
@@
Taking into account the caveats of our data set, which cannot discriminate between online and offline transactions, these results suggest that online marketplaces could also play a relevant role in explaining the increase in the number of low-value transactions in exports. They also underscore how decisions taken by few firms can have very large effects on the number of low-value trade transactions.
@@ 4 Conclusion
This paper has documented a novel feature of international trade: the increase in the share of low-value transactions in the volume of transactions. Using the universe of Spanish international trade transactions, we have shown that the share of low-value transactions in the total number of transactions increased from 9% to 61% in exports and from 14% to 54% in imports between 1997 and 2023.
We have offered two explanations for the increase in the number of low-value trade transactions. First, online retail platforms have allowed consumers to access a wide range of low-value items manufactured abroad, triggering a large increase in the volume of low-value transactions. Second, the fast-fashion strategy adopted by clothing retailers, which uses a frequent stocking of stores to adjust for changes in consumer preferences, has also led to a large increase in low-value transactions.
The increase in the share of low-value transactions can have important implications on international trade. Online platforms have allowed consumers, particularly those with low income, to have a more direct involvement in trade, providing them with a greater leverage in shaping trade flows. In addition, the increase in low-value import transactions may lead countries to use the de minimis threshold as a trade-restriction measure.
@@
Figure A.1: Robustness analyses on the share of low-value transactions in the total number of trade transactions
@@ A. Alternative low-value transaction thresholds
Note: Real de minimis corresponds to the 150 euro threshold, established in 2009, adjusted by the Spanish consumer price index (Spanish Statistical Institute, available at https://www.ine.es). US de minimis is the 800 USD de minimis value established by the United States in 2016. Benchmark corresponds to the share of low-value transactions reported in Panel A of Figure 1 in the main text.
@@ B. Top 20 countries by share in the total number of high-value transactions
Figure A.4: Ranking of destinations and origins by number of low-value transactions, 2021